# Intuitive Axial Augmentation Using Polar-Sine-Based Piecewise Distortion for Medical Slice-Wise Segmentation

### v1

@@ Abstract
Most data-driven models for medical image analysis rely on universal augmentations to improve performance. Experimental evidence has confirmed their effectiveness, but the unclear mechanism underlying them poses a barrier to the widespread acceptance and trust in such methods within the medical community. We revisit and acknowledge the unique characteristics of medical images apart from traditional digital images, and consequently, proposed a medical-specific augmentation algorithm that is more elastic and aligns well with radiology scan procedure. The method performs piecewise affine with sinusoidal distorted ray according to radius on polar coordinates, thus simulating uncertain postures of human lying flat on the scanning table. Our method could generate human visceral distribution without affecting the fundamental relative position on axial plane. Two non-adaptive algorithms, namely Meta-based Scan Table Removal and Similarity-Guided Parameter Search, are introduced to bolster robustness of our augmentation method. Experiments show our method improves accuracy across multiple famous segmentation frameworks without requiring more data samples. Our preview code is available in: https://github.com/MGAMZ/PSBPD.
@@
Medical Image Analysis , Image Augmentation , Piece-wise Affine , DICOM , Image Similarity , Image Preprocessing.
[…]
@@ 1 Introduction
Research Significance: Deep learning models have seen significant advancements in healthcare, including the widespread adoption of automated lesion segmentation, which alleviated the workload of radiologists and contributed to diagnostic precision. The up-to-date data-driven models often require a sufficient amount of data to obtain acceptable results, which can lead to unaffordable research costs in the medical field. In the computer vision community, various data augmentation techniques are widely used as part of preprocessing [2] to alleviate data scarcity. In the field of medical imaging, an excellent augmentation method can bring the following practical significance:
    Reduce the annotation cost of medical datasets.
    Fully tap the value of existing precious annotations without the need for more radiologists.
    Improve the robustness of automated annotation to enhance the “reliability” that is crucial in the medical field. Considering there’re plenty of existing neural networks for medical image analysis, we argue that creating a methodology that is applicable to most models is more impactful than solely focusing on developing an enhanced neural network model.
Traditional Augment: Effective but Counterintuitive. Recent research [9] suggests that some data augmentation methods transform the original data in more extreme ways (e.g. Erasing, Solarize) fig. 1. Medical experts often find these samples unreasonable, not present in the real world, and devoid of practical significance. However, the reality is not the case. Experimental results show that neural networks seem to still prefer these exaggerated enhancement methods, learning more knowledge from the “unreasonable” samples and showing better performance on downstream tasks. The reasons for this phenomenon are controversial, but it is clear that the interpretability of these data augmentation methods is very poor, which is particularly important in medical AI scenarios [28, 31].
Figure 1: The overview of our idea. (a) shows common frameworks tend to utilize universal augment method on CT dataset, without considering the effect of posture uncertainty. (b) is a pair of slices from the same subject with different posture, leading to variance on slices. Our augmentation method is able to simulate this malleable variance, thus generating more samples for neural networks.
@@
Our Method: Intuitive Distortion. According to the above, we proposed a Polar-sine-based Piecewise Affine Distortion specifically for medical radiologic image augmentation. Our approach, maps the reconstruction results of the original cross-sections to a distribution of which containing the subject’s posture changes. This is done to simulate the appearance when the body of the scanned individual exhibits slight distortions. In order to prevent disrupting the fundamental relative positional relationships of the human body’s tissues during augmentation, we construct a random transformation algorithm based on polar coordinates combined with sine functions. The scan table will be precisely identified and eliminated utilizing metadata-driven geometric algorithm. To align AI more closely with the intuitive aspects of medical practice, we incorporate similarity metrics to regulate the intensity of distortion, preventing unreasonable augmented samples. This approach also deliberately improves the efficiency of neural network search, thereby conserving resources during model fine-tuning. To our knowledge, this is the first data augmentation algorithm inspired by the differences in reconstructed slices due to uncertainties in human posture. Our preview code is available in: https://github.com/MGAMZ/PSBPD. In summary, our augmentation algorithm has the following features:
    It can generate any number of augmented sequences with differences from a single scan sequence, with controlable impact on the crucial relative positional features of human organs.
    Its hyperparameters are guided by similarity metrics, providing better interpretability and precision that are crucial in the medical field.
@@ 2 Related Works
Medical Radiologic Image Preprocess. [4] and [13] jumped out of the traditional idea of improving the accuracy of the model, and matched the physical imaging process of CT into the neural network, which greatly improved the contrast of images and the visibility of some difficult to observe tissues. This method is highly interpretable, but requires a calibration for each CT imaging device to obtain certain required parameters. This reduces its ease of use and accessibility. [30] emphasized the volumetric measurement of CT images and proposed a 2.5-D augmentation method. [15] skillfully isolated the image of the lesion area and then altered its background, which could even come from irrelevant samples. This approach aligns with deep learning practices and experiences [12, 5], yet it offers little in terms of explainability, which is crucial for medical applications. For instance, it is highly unconventional for images of the stomach and kidney to be present within the same slice.
Affine in Medical Image Processing. [8] reviewed the augmentation method to ease data scarcity of medical image. They pointed out that affine transformations (e.g., flip, rotation, translation, scaling, cropping and shearing) have widely used as a part of the pre-processing workflow for medical images. [9] further analyzed the latest research on augmentation and believed that augmentation is very effective for medical datasets. However, most studies have not been able to prove that their methods are effective for different models. [36] used multiple Affine Matrices on high-dimensional CT feature maps to differentially deform the vertebral bodies and surrounding soft tissues, leading to better registration performance. This is mainly due to the different elastic deformation characteristics between different tissues inside the body. [35] proposed an affine-enhanced arterial spin labeling (ASL) image registration method for MRI images. In this method, the affine transformation will be applied to image according to six parameters learned by deep learning neural network. [11] proposed a two-stage unsupervised learning framework for deformable medical image registration.
@@ 3.1 Data Preprocess Overview
Figure 2: Overview of the data preprocess. Symbol definitions are available in Table 1. The bold blue module is the proposed method. This approach is distinct and separate from neural networks and other processing, offering significant compatibility with various training frameworks.
@@ A Amplitude $[0,+\infty)$
Pixel Array Conversion will read reconstructed image stored in dcm file series using method proposed in [19]. Each patient’s scan includes multiple dcm files containing various meta data, which provides more possibilities for downstream tasks. The Resize operation is executed before the Distortion to achieve better performance, as the Distortion has an $O(H\times W)$ complexity. Normalization is performed after Distortion as a way to ensure that reconstructed pixel array conform to a standard distribution before being fed into the model.
@@ 3.2.1 Affine Control Point Initialization
To perform affine, we prepare a 2D grid of control points that are evenly distributed over the image. During the affine process, the pixels corresponding to the control points are moved along with the control points themselves. So, the density of control point map determines the affine accuracy, more points lead to more independent affine operations. We use the linspace function to generate an evenly spaced square grid with $\delta^{2}$ control points:
@@ 3.2.2 Affine Control Point Destination Calculation
Based on the previous description, we should ensure the following two points in the mapping transformation:
    The continuity relationship between image pixels remains unchanged.
    In order to conform to the distortion of the human body in reality, the distortion transformation should be reasonable comparing with actual scenarios.
@@
The key of the distortion is mapping each $\vartheta_{map}$ to a new location $\vartheta_{new}$. This conversion is performed with polar coordinate system, allowing us to easily control the absolute distance between each control point and the reconstruction center to remain constant, i.e. $(\frac{S_{h}}{\delta},\frac{S_{w}}{\delta})$. This satisfies the second requirement shown at the beginning of this chapter.
@@ 3.2.3 Piecewise Affine Execution
Now that the affine control point map $\mathcal{R}_{Smap}$ and its destination map $\mathcal{R}_{Tmap}$ has been generated, affine operation is available. As to piecewise affine, a Delaunay triangulation of the points is used to form a mesh ${\mathcal{R}_{\triangle}}^{N_{\triangle}\times 3\times 2}$ containing $i$ triangles. Delaunay triangulation function [7] is designed to maximize the minimum of all the angles of the triangles from a point set.
@@
$Delaunay$ tends to avoid narrow triangles, as these triangles can lead to extreme distortions in image transformations. One triangle ${R_{\triangle,i}}^{3\times 2}$ is composed of three control points $\left(\upsilon_{1},\upsilon_{2},\upsilon_{3}\right)\in\mathbb{Z}^{+}$, $i$ is the index of $\triangle$. The piecewise affine will apply customized transformations for each triangle. We assume $\mathcal{H}_{i}^{\triangle}$ as one triangle’s transformation matrix, and the piecewise affine can be described as (9).
@@ 3.3 Metadata-driven Scan Table Removal
In our research, the dataset comes from one hospital’s standardized DICOM data, which includes metadata [10] related to the scan table, thus allowing geometric modeling to remove the unneeded scanning bed’s HU value. The metadata and their symbols we use are illustrated in table 3. We define the physical table position as:
$\lambda=\frac{r}{\zeta}\times\varphi_{1}$
$For\quad p\quad in\quad\mathcal{R}_{pixels}:$
$\quad if\quad\left\lVert{p-\theta}^{2}\right\rVert>\lambda^{2}:$
$\quad\quad p=\varepsilon$
We give several examples of its effect in fig. 5 and fig. 7. The geometric definition is illustrated in fig. 6.
Figure 6: Geometric definitions used in scan table removal.
3.4 Similarity-Guided Hyperparameters Search
We employ SIFT (Scale-Invariant Feature Transform) [16, 17] and ORB (Oriented FAST and Rotated BRIEF) [27] to gauge the similarity of images pre- and post-augmentation [1]. These techniques efficiently highlight the feature discrepancy between images. Typically employed in image search, matching, and alignment [3, 6], we pioneer their application in the context of setting augmentation hyperparameters (i.e. $A,\omega$). Our improved training strategy with similarity guide is shown in algorithm 1.
Algorithm 1 Similarity-Guided Hyperparameters Search
  Similarity Guideline:
  Pretrain:
  $W_{pre}\leftarrow Pretrain(W_{0},D_{Train})$
  Fine-Tune:
  $W_{k}\leftarrow Train\left(W_{pre},D_{Train},\ k\right),\ \forall k\in K_{10\%}$
  $L_{k}\leftarrow Validation(W_{k},D_{Val},\ k),\ \forall k\in K_{10\%}$
  ${I_{A}}^{\ast},{I_{W}}^{\ast}=K^{\ast}\leftarrow{\arg\min}_{k\in K_{10\%}}L_{k}$
3.5 IRB Approval
The dataset is provided by Department of Gastrointestinal Surgery, Shanghai General Hospital. The dataset’s details are shown in 2 The hospital’s experts labeled the slice containing the largest gastric cancer area for each patient. The research was under the approval from Shanghai General Hospital Institutional Review Board (No. [2024] 032). The Approval Letter is available if required.
Item Describe
Source Shanghai General Hospital
Date Span 2014.12.26 $\sim$ 2021.09.18
Manufacturer GE MEDICAL SYSTEMS
Model Revolution CT
Slice Thickness 0.625mm / 1.25mm / 5mm
Software Versions sles_hde3.5 $\sim$ revo_ct_21b.32
Filter Body
4 Experiments
4.1 Experiment Basic Settings
The invalid area in dcm pixel array is represented by a negative constant value of $-2000$. These padding values significantly affect the data value distribution, skewing the mean, and reducing the fitting speed fig. 8. All dcm metadata used in this research are listed in table 3, patients’ private information is not included.
Additionally, we utilized publicly available large-scale datasets to validate the effectiveness of our proposed method, i.e. AbdomenCT1K[18, 33] and CT-ORG[25, 33].
Figure 8: Pixel value distribution. The number of $-2000$ values is way far from the those of the other values, which correspond to padding values. So, we zoomed in values larger than zero to filter padding points. The major valid values are concentrated in two ranges, between $0\sim 200$ (1) and $850\sim 1200$ (2). Noted that (3) is of log scale. In general, negative effective values are less common and tend to cluster around the minimum value of $-1024$.
@@ 4.1 Experiment Basic Settings
Reconstruction Diameter 3 $S/2$ G 0018 E 1100 DS Diameter, in mm, of the region from within which the data was used in creating the reconstruction of the image.

### v2

# mode = titleAxial Distortion

@@ Abstract
Most data-driven models for medical image analysis rely on universal augmentations to improve accuracy. Experimental evidence has confirmed their effectiveness, but the unclear mechanism underlying them poses a barrier to the widespread acceptance and trust in such methods within the medical community. We revisit and acknowledge the unique characteristics of medical images apart from traditional digital images, and consequently, proposed a medical-specific augmentation algorithm that is more elastic and aligns well with radiology scan procedure. The method performs piecewise affine with sinusoidal distorted ray according to radius on polar coordinates, thus simulating uncertain postures of human lying flat on the scanning table. Our method could generate human visceral distribution without affecting the fundamental relative position on axial plane. Two non-adaptive algorithms, namely Meta-based Scan Table Removal and Similarity-Guided Parameter Search, are introduced to bolster robustness of our augmentation method. In contrast to other methodologies, our method is highlighted for its intuitive design and ease of understanding for medical professionals, thereby enhancing its applicability in clinical scenarios. Experiments show our method improves accuracy with two modality across multiple famous segmentation frameworks without requiring more data samples. Our preview code is available in: https://github.com/MGAMZ/PSBPD.
@@ 1 Introduction
Research Significance. Deep learning models have seen significant advancements in healthcare, including the widespread adoption of automated lesion segmentation, which alleviated the workload of radiologists and contributed to diagnostic precision. The up-to-date data-driven models often require a sufficient amount of data to obtain acceptable results, which can lead to unaffordable research costs in the medical field. In the computer vision community, various data augmentation techniques are widely used as part of preprocessing [3] to alleviate data scarcity. In the field of medical imaging, an excellent augmentation method should meet several requirements: 1) Reduce the annotation cost of medical datasets. 2) Fully tap the value of existing precious annotations without the need for more radiologists. 3) Being easily comprehensible to doctors in clinical scenarios, thus ensuring practical applicability [28, 13]. Considering there’re plenty of existing neural networks for medical image analysis, we argue that creating a methodology that is applicable to most models is more impactful than solely focusing on developing an enhanced neural network model.
Traditional Augment: Effective but Counterintuitive. Recent research [10] suggests that some data augmentation methods transform the original data in more extreme ways (e.g. Erasing, Solarize, cross-domain synthesis, and generative models, etc..). Medical experts often find these samples unreasonable, not present in the real world, and devoid of practical significance. Samples generated through these techniques may not conform strictly to anatomical standards. For example, in a CT scan sequence, if the HU value for the heart region is set to a constant and the locations of the bladder and liver are interchanged, such samples are generally classified as noise and are removed during the standard data analysis procedures [17, 34]. However, the reality is not the case as to deep learning. Experimental results show that neural networks seem to still prefer these exaggerated enhancement methods, learning more knowledge from the “unreasonable” samples and showing better accuracy on downstream tasks. The reasons for this phenomenon are controversial, but it is clear that the interpretability of these data augmentation methods is very poor, which is particularly important in medical AI scenarios [32, 36].
Physicians typically scrutinize how data is collected and its quality. If they find out a model has been trained on deliberate “noise” data and claims a promising outcome, they’re likely to be skeptical. It’s counterintuitive to deliberately introduce noise for better accuracy, particularly when we can’t clearly justify “why it can.” As a result, enhancing the level of interpretability and optimizing the system’s transparency contribute to the potential application of a method in clinical scenarios.
@@
Figure 1: Overview of the proposed method.
(a) shows common frameworks tend to utilize universal augment method on CT dataset, without considering the effect of posture uncertainty. (b) is a pair of slices from the same subject with different posture, leading to variance on slices. Our augmentation method is able to simulate this malleable variance, thus generating more samples for neural networks.
As previously noted, simulation algorithms developed in response to this observation offer enhanced interpretability. In clinical practice, it is expected that the same patient will yield varying reconstruction results across different scans. Consequently, the samples produced by these methods are deemed to align with clinical realities and do not introduce any inexplicable elements. It is a well-acknowledged fact within the radiology community that radiologists often encounter studies that include multiple series, with each series displaying similar yet subtly different human body postures [20, 1]. Medical professionals will readily comprehend the enhancement in accuracy achieved through the use of these samples. Because this is analogous to gathering additional data on a patient, which inherently aids in the intuitive understanding and facilitation of any data analysis process.
Our Method: Intuitive Distortion. According to the above, we proposed a Polar-sine-based Piecewise Affine Distortion specifically for medical radiologic image augmentation. Our approach, maps the reconstruction results of the original cross-sections to a distribution of which containing the subject’s posture changes. This is done to simulate the appearance when the body of the scanned individual exhibits slight distortions. In order to prevent disrupting the fundamental relative positional relationships of the human body’s tissues during augmentation, we construct a random transformation algorithm based on polar coordinates combined with sine functions. The scan table will be precisely identified and eliminated utilizing metadata-driven geometric algorithm. The elimination helps to the undesirable artifacts introduced by the augmentation. To align AI more closely with the intuitive aspects of medical practice and help reduce the cost when applying the augmentation to a new dataset, we incorporate similarity metrics to regulate the intensity of distortion, preventing unreasonable augmented samples. This Non-Neural approach also improves the efficiency of neural network search, thereby conserving resources during model fine-tuning.
In order to validate our methodology, we assembled a high-quality dataset of gastric cancer CT scans, which are collected over a period of seven years during clinical practice. This dataset is meticulously annotated by two medical professionals, resulting in the creation of a few-anno CT dataset comprising 689 studies. In order to ensure the reproducibility and compatibility of our proposed method, we incorporated two public datasets and train from scratch on seven distinct models across two paradigms. The empirical evidence indicates that our method not only maintains a satisfactory level of interpretability but also improve the precision across the majority of applicable scenarios.
Our preview code is available in: https://github.com/MGAMZ/PSBPD. In summary, our augmentation algorithm has the following features:
    It can generate any number of augmented sequences with differences from a single scan sequence, with controlable impact on the crucial relative positional features of human organs to align with anatomical intuition, thus increasing the potential of applications within clinical scenarios.
    The application of similarity metrics facilitates a swift determination of the optimal parameter spectrum, thereby augmenting the deployment efficacy on diverse datasets.
@@ 2 Related Works
Medical Radiologic Image Preprocess. [5] and [15] jumped out of the traditional idea of improving the accuracy of the model, and matched the physical imaging process of CT into the neural network, which greatly improved the contrast of images and the visibility of some difficult to observe tissues. This method is highly interpretable, but requires a calibration for each CT imaging device to obtain certain required parameters. This reduces its ease of use and accessibility. [35] emphasized the volumetric measurement of CT images and proposed a 2.5-D augmentation method. [17] skillfully isolated the image of the lesion area and then altered its background, which could even come from irrelevant samples. This approach aligns with deep learning practices and experiences [14, 6], yet it offers little in terms of explainability, which is crucial for medical applications. For instance, it is highly unconventional for images of the stomach and kidney to be present within the same slice. The approach we are suggesting has no need for hardware calibration. Although certain procedures do require DICOM metadata, these are not mandatory.
Affine in Medical Image Processing. [9] reviewed the augmentation method to ease data scarcity of medical image. They pointed out that affine transformations (e.g., flip, rotation, translation, scaling, cropping and shearing) have widely used as a part of the pre-processing workflow for medical images. [10] further analyzed the latest research on augmentation and believed that augmentation is very effective for medical datasets. [43] used multiple Affine Matrices on high-dimensional CT feature maps to differentially deform the vertebral bodies and surrounding soft tissues, leading to better registration accuracy. This is mainly due to the different elastic deformation characteristics between different tissues inside the body. The method aligns with our thinking in assuming the human body’s non-rigid nature. However, our mapping is more granular and extends beyond just local tissue structures. [42] proposed an affine-enhanced arterial spin labeling (ASL) image registration method for MRI images. In this method, the affine transformation will be applied to image according to six parameters learned by deep learning neural network. [12] proposed a two-stage unsupervised learning framework for deformable medical image registration. This method features a larger granularity and integrates a complicated two-phase modeling strategy.
@@ 3.1 Data Preprocess Overview
Figure 2: Overview of data preprocess.
Symbol definitions are available in Table 1. The bold blue module is the proposed method. This approach is distinct and separate from neural networks and other processing, offering significant compatibility with various training frameworks. In the context of data flow, the proposed method is positioned subsequent to size normalization and prior to normalization. The inclusion of Scan Table Removal is optional, serving to alleviate potential negative effects resulting from the proposed augmentation. Utilization of Similarity Metrics aids in estimating viable $A,\omega$, thereby enhancing the efficiency of parameter search.
@@ A Amplitude $[0,+\infty)$
Pixel Array Conversion will read reconstructed image stored in dcm file series using method proposed in [21]. Each patient’s scan includes multiple dcm files containing various meta data, which provides more possibilities for downstream tasks. The Resize operation is executed before the Distortion to achieve better accuracy, as the Distortion has an $O(H\times W)$ complexity. Normalization is performed after Distortion as a way to ensure that reconstructed pixel array conform to a standard distribution before being fed into the model.
@@ 3.2.1 Affine Control Point Initialization
To perform affine, we prepare a 2D grid of control points that are evenly distributed over the image. Control points function by establishing the mapping for a select few coordinates, thereby directing the transformation of the entire 2D matrix. The pixels corresponding to these control points remain constant before and after the mapping process, and they move in unison to the new location, adhering to the principles of classical piecewise affine mapping. During the affine process, the pixels corresponding to the control points are moved along with the control points themselves. So, the density of control point map determines the affine accuracy, more points lead to more independent affine operations. We use the linspace function to generate an evenly spaced square grid with $\delta^{2}$ control points:
@@ 3.2.2 Affine Control Point Destination Calculation
Based on the previous description, we should ensure the following two points in the mapping transformation: 1) The continuity relationship between image pixels remains unchanged, and 2) the distortion transformation should be reasonable comparing with actual scenarios, also conform to the distortion of the human body in reality.
@@
The key of the distortion is mapping each $\vartheta_{map}$ to a new location $\vartheta_{new}$. This conversion is performed with polar coordinate system, allowing us to easily control the absolute distance between each control point and the reconstruction center to remain constant, i.e. $(\frac{S_{h}}{\delta},\frac{S_{w}}{\delta})$. This satisfies the second requirement shown at the beginning of this chapter. In actual scenario, human body’s distortion amplifies where is close to the body’s surface. In other words, a positive correlation between the distortion and the distortion center. When $\omega$ is not excessively large, the mapping of polar angles for all control points associated with a single polar angle tends to be comparable. Conversely, when W is extremely close to zero, the mapping resembles a rotational transformation, which barely impacts regional features.
@@ 3.2.3 Piecewise Affine Execution
Now that the affine control point map $\mathcal{R}_{Smap}$ and its destination map $\mathcal{R}_{Tmap}$ has been generated. Next, it is necessary to derive the pixel-level sampling relationship based on the mapping relationship of these control points, in order to sample a new matrix from the source matrix, which also means obtaining new samples. As to piecewise affine, a Delaunay triangulation of the points is used to form a mesh ${\mathcal{R}_{\triangle}}^{N_{\triangle}\times 3\times 2}$ containing $i$ triangles. Delaunay triangulation function [8] is designed to maximize the minimum of all the angles of the triangles from a point set.
@@
$Delaunay$ tends to avoid narrow triangles, as these triangles can lead to extreme distortions in image transformations. One triangle ${R_{\triangle,i}}^{3\times 2}$ is composed of three control points $\left(\upsilon_{1},\upsilon_{2},\upsilon_{3}\right)\in\mathbb{Z}^{+}$, $i$ is the index of $\triangle$. The piecewise affine will apply customized transformations for each triangle. We assume $\mathcal{H}_{i}^{\triangle}$ as one triangle’s transformation matrix, and the piecewise affine can be described as Eq. 9.
@@ 3.3 Metadata-driven Scan Table Removal
In the usual course of CT reconstruction, the resulting sequences often include images of the CT Scan Table, predominantly positioned directly beneath the patient. Despite variations among subjects, the position of the CT Scan Table remains relatively constant. Consequently, the neural network, upon encountering this structure frequently in a large sample set and being informed by annotations that it is not a target of interest, tends to disregard the CT Scan Table and rarely misclassifies it as a relevant target. Fig. 4 illustrates not only the slice morphology under varying augmentation intensities but also highlights a concerning issue. The proposed augmentation method, which applies affine mapping to all pixels across the entire slice, transforms the CT Scan Table into a complex, multi-segment curve shape. Additionally, minor discrepancies in the Table’s position in the original space are exaggerated during the augmentation process, which can be interpreted as noise enhancement and is considered undesirable.
The crux of this step lies in accurately identifying the pixels in each slice that correspond to the CT Scan Table. While a two-stage machine learning model that incorporates identification followed by segmentation is a viable approach, it entails a substantial increase in computational complexity. Capitalizing on the inherent benefits of medical sequence imaging, we employ DICOM metadata to spatially model the CT sequence, thereby precisely determining the location of the CT Scan Table. This method eschews the use of adaptive modules, thereby preventing an escalation in computational overhead. As a preparation, we deliberately preserved the DICOM metadata (while ensuring necessary anonymization) [11]. All metadata and their symbols we use are illustrated in Table 2. The overview of geometric positional modeling is illustrated in Fig. 5.
@@
Reconstruction Diameter 3 $S/2$ G 0018 E 1100 DS Diameter, in mm, of the region from within which the data is used in creating the reconstruction of the image.
@@
Figure 5: Geometric Positional Modeling used in scan table removal. The definition of DICOM symbols are available in Table 2.
Among the points in Fig. 5, we define the physical table position as:
$\lambda=\frac{r}{\zeta}\times\varphi_{1}$
$For\quad p\quad in\quad\mathcal{R}_{pixels}:$
$\quad if\quad\left\lVert{p-\theta}^{2}\right\rVert>\lambda^{2}:$
$\quad\quad p=\varepsilon$
It is important to note that during the similarity detection process, our objective is to ascertain whether the sample has been excessively distorted. This process is independent of the neural network and maintains an equitable treatment of all pixels, thus the CT Scan Table Removal feature remains inactive during the detection phase. We give several examples of its effect in Fig. 6 and Fig. 7.
The CT Scan Table Removal module necessitates specific DICOM metadata or analogous parameters for its operation, which are generally not available in most public datasets (formatted as NIFTI). Consequently, the application of this module is constrained. Within the scope of our proposed research, this module is not deemed essential. Its utilization is anticipated to enhance the accuracy of subsequent neural networks; however, the accuracy expectations should remain satisfactory even in its absence.
3.4 Similarity-Guided Hyperparameters Search
The proposed distortion involves two key parameters (i.e. $A,\omega$) that are specifically designed to control the degree of augmentation. Increasing the value of parameter $A$ results in a higher degree of distortion. Similarly, elevating the value of parameter $\omega$ causes the proposed distortion mapping to resemble a rotational affine more closely. While the intensity of the distortion can be intuitively assessed through visualization, this method does not provide a basis for estimating the accuracy of end-to-end predictions. Hence, we proceed from the assumption that excessive augmentation leads to the destruction of image features, which hinders the neural network’s ability to discern effective features, thereby complicating the training process. We employ SIFT (Scale-Invariant Feature Transform) [18, 19] and ORB (Oriented FAST and Rotated BRIEF) [31] to gauge the similarity of images pre- and post-augmentation [2]. At lower augmentation intensities, the sample undergoes minimal changes, allowing the similarity algorithm to identify a greater number of corresponding points between the pre- and post-augmentation samples. However, as the features become compromised, the number of corresponding points diminishes. The performance of similarity detection significantly exceeds that of deep neural networks and remains acceptable when executed on a CPU.
This indicates that we can utilize these metrics to quantify the level of feature degradation. An effective augmentation technique should strike a balance between enhancing the dataset’s information entropy and preserving the learnability of the samples for the neural network. Typically employed in image search, matching, and alignment [4, 7], we pioneer their application in the context of setting augmentation hyperparameters (i.e. $A,\omega$). Our improved training strategy with similarity guide is shown in Algorithm 1. Under the auspices of the Similarity Guide, the FineTune process is streamlined to conduct its search within the $K_{10\%}$ parameter spectrum.
Algorithm 1 Similarity-Guided Hyperparameters Search
Similarity Guideline:
Pretrain:
$W_{pre}\leftarrow Pretrain(W_{0},D_{Train})$
Fine-Tune:
$W_{k}\leftarrow Train\left(W_{pre},D_{Train},\ k\right),\ \forall k\in K_{10\%}$triangleright With Distortion and Table Removal
$L_{k}\leftarrow Validation(W_{k},D_{Val},\ k),\ \forall k\in K_{10\%}$triangleright With Table Removal
${I_{A}}^{\ast},{I_{W}}^{\ast}=K^{\ast}\leftarrow{\arg\min}_{k\in K_{10\%}}L_{k}$
3.5 Efficiency
The components of the method we have developed are designed to function independently of the neural network. These components can be executed concurrently with neural network computations during the data preprocessing stage, enhancing overall efficiency. Initializing the control point matrix $\mathcal{R}_{Smap}$ has a space complexity of $O(\delta^{2})$, and it merely requires memory allocation with almost no additional time cost. Each sample necessitates an individual computation for every control point. The $\mathcal{R}_{Tmap}$ is derived by iterating over all control points. The time complexity is linearly dependent on the number of control points, expressed as $O(\delta^{2})$. The computation for each control point entails two conversions between Cartesian and polar coordinate systems, in addition to a polar angle mapping procedure. These operations can be efficiently processed by CPUs.
For triangulation, the Bowyer-Watson based Delaunay method [27] is a prominent technique, exhibiting an average time complexity of $O(nlogn)$. In this context, $n$ represents the total count of pixel points. Given that CT cross-sections commonly feature an initial reconstruction of $512\times 512$ pixels, this step is recognized as a computationally intensive, pointwise operation. Next, affine mapping point calculations need to be performed for each sub-triangle. We use the affine estimation method provided by Scipy [37], and its maximum computational complexity arises from the SVD calculation, which is $O(\delta^{3})$. Upon the completion of this step, we will have established the point-to-point mapping matrix. For each pixel in the output matrix $\mathcal{R}_{distorted}$ , the corresponding source pixel coordinates are retrieved from the mapping matrix (Eq. 9), and sampling is conducted with these coordinates as the center from the source matrix. To mitigate distortion artifacts, bicubic interpolation is utilized during the sampling process. This interpolation technique exhibits a computational complexity of $O(n^{2})$.
In summary, the computational overhead of the method we have proposed is primarily associated with the determination of the affine mapping matrix $\mathcal{H}$ and the piecewise affine mapping sampling process. By reducing the resolution $\delta$ of the control point matrix, we can decrease the computational time required for the affine mapping matrix. Similarly, lowering the resolution $\mathcal{S}_{h}\times\mathcal{S}_{w}$ of the source matrix can reduce the duration of the sampling execution.
3.6 IRB Approval
The dataset is provided by Department of Gastrointestinal Surgery, Shanghai General Hospital. The dataset’s details are shown in Table 3 The hospital’s experts labeled the slice containing the largest gastric cancer area for each patient. The research is under the approval from Shanghai General Hospital Institutional Review Board (No. [2024] 032). The Approval Letter is available if required.
Item Describe
Source Shanghai General Hospital
Date Span 2014.12.26 $\sim$ 2021.09.18
Manufacturer GE MEDICAL SYSTEMS
Model Revolution CT
Slice Thickness 0.625mm / 1.25mm / 5mm
Software Versions sles_hde3.5 $\sim$ revo_ct_21b.32
Filter Body
4 Experiments
4.1 Data Collection
In order to validate the efficacy of our augmentation method, we compiled CT scan sequences from 895 patients diagnosed with gastric cancer through clinical practice. A gastrointestinal surgeon identified the slice with the largest gastric cancer lesion area for each patient’s scan, while a radiologist provided detailed cancerous region annotations on this slice, establishing a pixel-level segmentation benchmark. The gastrointestinal surgeon conducted a review of the annotations and made necessary revisions. This approach ensures that only one slice is annotated regardless of the total number of slices in the CT scan sequence. Through a collaborative effort, the radiologist and gastrointestinal surgeon achieved consensus on all scan sequences and segmentations. Following a cleaning process, which excluded cases with 1) severe artifacts, 2) unsatisfactory imaging, 3) unavailable pathological results, 4) incomplete lesion coverage, 5)indeterminable maximum cross-sectional area due to small lesion size, 6) patient or family objections, and 7) unavailability of DICOM metadata, we selected 689 cases from the initial 895 for further research.
The dataset incorporated DICOM files with standardized metadata, while the annotation data is preserved separately in nrrd format. The nrrd files contained the original section coordinates, which are utilized to correlate with the DICOM sequences and ascertain the specific section for each annotation. The 895 image sequences collected encompassed reconstruction layer thicknesses of 0.625mm, 1.25mm, and 5mm, with no uniform standard for slice spacing. To ensure consistency, all sequences are resampled to a voxel size of $1mm\times 1mm\times 1mm$, facilitating subsequent training processes.
Additionally, we utilized publicly available large-scale datasets to validate the effectiveness of our proposed method, BraTS[22] and CT-ORG[29, 40].
4.2 Metrics and Baselines
We use three pixel-level metrics, Dice, Recall, and Precision, to determine the model’s accuracy in identifying lesions or tissues. The computation formulas for the three metrics are presented below.
$Dice=\frac{2\times TP}{2\times TP+FP+FN}$ (13)
$Recall=\frac{TP}{TP+FN}$
$Precision=\frac{TP}{TP+FP}$
where $TP$ represents the number of true positive pixels, $FP$ represents the number of false positive pixels, and $FN$ represents the number of false negative pixels.
Among the techniques employed for data augmentation in interpretable medical image volume sequences, rotation enhancement stands out as one of the most extensively applied and reliable methods [39]. In terms of neural networks, there are currently two main architechtures, which are the convolutional structure and the Transformer structure. For the public dataset, we have incorporated the SwinUMamba model, which exhibits significant potential, to assess the sensitivity of our method to state space models. At present, the DICOM metadata for most commonly utilized public datasets is not openly available, which constrains the thoroughness of evaluation processes.